OpenStudy (anonymous):

Transform10sqrt(10) + 11sqrt(11) into the equivalent exact answer: sqrt( 2331+220sqrt(110) )

OpenStudy (anonymous):

\[10\sqrt{10}+11\sqrt{11}=\sqrt{2331+220\sqrt{110}}\]

OpenStudy (anonymous):

I want to know how to get the other answer

OpenStudy (welshfella):

expand (10sqrt10 + 11sqrt11)^2 first then take square root

OpenStudy (anonymous):

thank you